# SpectrumFM: A Foundation Model for Intelligent Spectrum Management

### v1

@@ Abstract
Index Terms:
Spectrum foundation model, automatic modulation classification, wireless technology classification, spectrum sensing, anomaly detection.
@@ I Introduction
The radio frequency (RF) spectrum is a fundamental yet increasingly scarce resource in modern wireless communication networks [1]. As the demand for ubiquitous connectivity continues to rise, driven by the proliferation of Internet-of-Things (IoT) devices, non-terrestrial networks (NTNs), and the sixth-generation (6G) wireless paradigms, spectrum management is of crucial importance for improving spectrum efficiency, mitigating interference, detecting the illegal spectrum utilization, achieving reliable and secure spectrum operation [2, 3].
@@
Inspired by the transformative success of foundation models in natural language processing (e.g., BERT, GPT) and computer vision (e.g., ViT, Segment Anything), there is growing interest in developing foundation models for wireless communications [8, 9]. Those models are trained on large-scale heterogeneous data to learn universal representations that can be adapted to diverse downstream tasks with minimal supervision. Recent efforts, such as SpectralGPT [10] and WirelessGPT [11], have demonstrated the potential of this paradigm in remote sensing and integrated sensing and communication (ISAC), respectively. However, the spectrum environment is inherently dynamic, with interference, signal fading, and noise introducing significant variability. This variability complicates the model ability to generalize across different environmental conditions and spectrum usage patterns. Moreover, the spectrum encompasses a broad range of frequency bands, from licensed to unlicensed, and the unpredictable nature of cognitive radio systems further amplifies the challenges. As a result, the application of foundation models to spectrum management remains largely unexplored.
@@ II-A Task-Specific Spectrum Management Methods
Spectrum management has traditionally been approached through task-specific methods, where models are designed and optimized for individual tasks such as AMC, WTC, SS, and AD. AMC, in particular, is a challenging task that requires accurate classification of the modulation scheme from a given IQ signal. Traditional methods have primarily relied on hand-crafted features, such as cyclic-moment-based features extracted through expert domain knowledge [12]. A significant milestone was achieved in [13], where CNNs were introduced to learn features automatically, eliminating the dependence on expert-designed representations. This marked the advent of deep learning techniques in AMC.
@@
WTC is another challenging task that requires the model to accurately classify the type of wireless transmission based on the input IQ signal. Numerous studies have demonstrated significant progress in this area. For example, the authors in [19] employed CNNs to automatically extract features from the input signal and used a feedforward neural network to classify the wireless technology. Data augmentation techniques were also incorporated, enabling the model to generalize well to unseen scenarios while maintaining high accuracy. In [20], a multiscale module was introduced to capture features at various resolutions. Additionally, a global average pooling layer was utilized to aggregate the extracted features, which reduced the number of parameters and accelerated the training process. Despite its simplified architecture, the model achieved strong classification performance and fast convergence.
The authors in [21] proposed a comprehensive WTC system designed to identify and analyze a wide range of wireless technologies coexisting within the 5.9 GHz band. The system specifically focused on LTE, Wi-Fi, 5G NR, C-V2X PC5, and ITS-G5. To ensure accurate identification, the study adopted a short time resolution window based on the minimum frame duration of the target technologies. Unlike most IQ-based approaches, MFBLN [22] leveraged alternative features such as received signal strength indicator (RSSI), fast fourier transform (FFT), and spectrogram representations to extract multi-dimensional signal characteristics. This allowed the model to capture richer information from the input, thereby enhancing classification accuracy and robustness across diverse wireless environments.
SS is a fundamental task in cognitive radio networks, referring to the process of monitoring and analyzing the RF spectrum to determine its occupancy status [23]. The primary goal of SS is to identify spectrum holes, which are unused or underutilized frequency bands in a specific geographical area. By detecting these holes, cognitive radios can dynamically access the available spectrum without interfering with primary users (PUs). SS is generally categorized into two types, namely narrowband SS and wideband SS. Narrowband SS analyzes one frequency band at a time, while wideband SS simultaneously evaluates multiple bands to identify vacant channels suitable for transmission [24]. Common narrowband SS techniques include energy detection, cyclostationary detection, and matched filter detection [25]. Among these, energy detection is widely used due to its simplicity. It determines spectrum occupancy by comparing the measured signal energy against a predefined threshold.
@@ II-B Foundation Models
Foundation models are large-scale machine learning models that are pre-trained on vast amounts of data to learn general-purpose representations. The models serve as a foundational basis for a wide range of downstream tasks, making them highly versatile and adaptable across various domains. Foundation models are typically developed using deep learning techniques and are characterized by their massive scale, extensive training data, and strong transferability [33, 34]. The authors in [35] introduced the attention mechanism and the Transformer architecture, which serve as the cornerstone of foundation models. BERT [36] pioneered the era of foundation models. By introducing a bidirectional Transformer-based architecture and innovative pre-training strategies such as masked language modeling (MLM) and next sentence prediction (NSP), BERT achieved state-of-the-art performance on a wide range of nature language processing (NLP) tasks.
The BERT-like prediction loss is employed in HuBERT [37] for self-supervised speech representation learning, leading to the development of a speech foundation model. The authors in [38] present a foundational model that reformulates and unifies four core vision tasks, namely object detection, instance segmentation, keypoint prediction, and image captioning, into a single pixel-to-sequence framework. By framing multiple vision tasks within a shared pixel-to-sequence paradigm, this work demonstrated the potential of foundation models to generalize across tasks, reducing the need for specialized architectures and promoting greater flexibility in multimodal applications. SpectralGPT [10] represents a pioneering effort in creating a universal foundation model tailored for remote sensing (RS) applications. Trained on a large-scale dataset comprising one million spectral RS images, SpectralGPT achieved a parameter size exceeding 600 million. Evaluations across four downstream tasks reveal substantial performance improvements, highlighting the potential of foundation models to tackle a wide range of RS tasks.
WirelessGPT [11], a pioneering foundation model tailored for multi-task learning in ISAC, features an initial parameter size of approximately 80 million. Pre-trained on a large-scale wireless channel dataset, it delivers significant advancements in critical wireless tasks such as channel estimation, channel prediction, human activity recognition, and wireless reconstruction. WirelessGPT demonstrates that foundation models can achieve impressive performance improvements over conventional methods and smaller artificial intelligence (AI) models while requiring fewer labeled data, thereby enhancing efficiency and scalability in ISAC applications.
@@ III Our Proposed Model
$N$ The number of signal symbols.
@@ III-A Data Collection And Processing Stage
In the data collection and processing stage, a diverse set of spectrum data is gathered from multiple sources, including publicly available open-source datasets and signals collected in real-world for this study, to facilitate comprehensive training of SpectrumFM. The open-source datasets utilized in this study include the RML2018.01A dataset [39] and the technology recognition (TechRec) dataset [^2] https://ieee-dataport.org/documents/iq-signals-captured-lte-wifi-and-dvb-t. The RML2018.01A dataset comprises 24 distinct modulation types, including OOK, 4-ASK, 8-ASK, BPSK, QPSK, 8-PSK, 16-PSK, 32-PSK, 16-APSK, 32-APSK, 64-APSK, 128-APSK, 16-QAM, 32-QAM, 64-QAM, 128-QAM, 256-QAM, AM-SSB-WC, AM-SSB-SC, AM-DSB-WC, AM-DSB-SC, FM, GMSK, and OQPSK. This dataset spans a broad range of SNRs, ranging from -20 dB to 30 dB in increments of 2 dB, providing a comprehensive representation of diverse signal conditions. The TechRec dataset consists of over-the-air IQ samples collected from six distinct locations in Gent, Belgium, including UZ, Reep, Rabot, Merelbeke, iGent, and Gentbrugge. Each location exhibits unique signal characteristics due to variations in signal strength, noise levels, and physical propagation conditions. The SNR in the dataset ranges from -6 dB to 12 dB, with increments of 2 dB. Specifically, IQ samples collected from Merelbeke, iGent, and Rabot are leveraged for pre-training, while data from UZ, Reep, and Gentbrugge are exploited for fine-tuning and testing. In addition to the publicly available datasets, signals are also collected by using a ceyear 1435B-V RF signal generator for transmission and a SAM 60 MK2 receiver for reception. The total size of the pre-training dataset amounts to approximately 25 GB. The received signal $H$ is formulated as
@@
To facilitate normalization, we first convert the IQ signals into amplitude-phase (AP) signals, since the value ranges of amplitude and phase components are better suited for standardization, specifically $[-\pi,\pi]$ for the phase component and $[0,+\infty]$ for the amplitude component. Mathematically, this transformation for the IQ signals to AP signals can be expressed as
@@ III-B SpectrumFM Encoder
where $\mathbf{x}_{\text{proj}}\in\mathbb{R}^{N\times d}$ is the projected signal, $\mathbf{W}_{\text{proj}}\in\mathbb{R}^{d\times 2\times 1}$ is the convolution kernel, $d$ is the latent dimension, and $p$ represents the position index, ranging from 0 to $N-1$. To capture the sequential order in a signal sequence, positional encoding is added to $\mathbf{x}_{\text{proj}}$. The positional encoding is given as
@@
where $i$ is the index of the dimension, ranging from 0 to $d-1$. The signal with the positional encoding is given as
@@
where concat()($\cdot$)( ⋅ ) is the concatenation operator, and $\mathbf{W}^{o}\in\mathbb{R}^{d\times d}$ is the weight matrix.
@@ III-C1 The Masked Reconstruction Task
In the masked reconstruction task, the objective is to mask a portion of the original AP sequence $\mathbf{x}_{\text{ap}}^{\text{norm}}$ and train the model to reconstruct the masked parts. The task facilitates to learn general representations in an unsupervised manner by training the model to infer and reconstruct masked portions of the input data, thereby enhancing its robustness and generalization capabilities. Specifically, given the sequence $\mathbf{x}_{\text{ap}}^{\text{norm}}$, we first randomly mask a subset of the sequence according to a predefined masking ratio $r$. The binary mask vector $\mathbf{m}$ is defined as $\mathbf{m}\sim\text{Bernoulli}(1-r)$, which consists of elements $\left\{m_{1},m_{2},\cdots,m_{N}\right\}$. Here $m_{p}=0$ if the $p$-th symbol is masked out, and $m_{p}=1$ otherwise. The masked sequence is then given as
@@
where $\mathbf{M}$ assigns extremely small values (e.g., $\infty$- ∞) to the positions that are masked out, thereby ensuring that the positions do not contribute to the computed attention weights. After obtaining $\mathbf{x}_{\text{{latent}}}^{(L)}$, a lightweight decoder is applied to the latent representation to reconstruct the original input sequence, given as
@@ III-C2 Next-slot Signal Prediction Task
The next-slot signal prediction task aims to predict the signal symbol at the $N$-th time slot based on the historical signal observations from the previous $N-1$ slots. The task facilitates SpectrumFM to understand the temporal characteristics of signals. Specifically, $\mathbf{x}_{\text{ap}}^{\text{norm}}[:N-1]$ is input to the SpectrumFM encoders to obtain the latent representation $\mathbf{x}_{\text{{latent}}}^{(L)}$, and then an long short-term memory (LSTM) layer is leveraged to aggregate the latent representation, and finally a linear layer is exploited to predict the $N$-th signal symbol. Mathematically, the predicted signal symbol is given as

### v2

@@ I Introduction
The radio frequency (RF) spectrum is a fundamental yet increasingly scarce resource in modern wireless communication networks [meaning]. As the demand for ubiquitous connectivity continues to rise, driven by the proliferation of Internet-of-Things (IoT) devices, non-terrestrial networks (NTNs), and the sixth-generation (6G) wireless paradigms, spectrum management is of crucial importance for improving spectrum efficiency, mitigating interference, detecting the illegal spectrum utilization, achieving reliable and secure spectrum operation [10488747, 10685064].
@@
Inspired by the transformative success of foundation models in natural language processing (e.g., BERT, GPT) and computer vision (e.g., ViT, Segment Anything), there is growing interest in developing foundation models for wireless communications [10876763, 10841938]. Those models are trained on large-scale heterogeneous data to learn universal representations that can be adapted to diverse downstream tasks with minimal supervision. Recent efforts, such as SpectralGPT [10490262] and WirelessGPT [yang2025wirelessgptgenerativepretrainedmultitask], have demonstrated the potential of this paradigm in remote sensing and integrated sensing and communication (ISAC), respectively. However, the spectrum environment is inherently dynamic, with interference, signal fading, and noise introducing significant variability. This variability complicates the model ability to generalize across different environmental conditions and spectrum usage patterns. Moreover, the spectrum encompasses a broad range of frequency bands, from licensed to unlicensed, and the unpredictable nature of cognitive radio systems further amplifies the challenges. As a result, the application of foundation models to spectrum management remains largely unexplored.
@@ II-A Task-Specific Spectrum Management Methods
Spectrum management has traditionally been approached through task-specific methods, where models are designed and optimized for individual tasks such as AMC, WTC, SS, and AD. AMC, in particular, is a challenging task that requires accurate classification of the modulation scheme from a given IQ signal. Traditional methods have primarily relied on hand-crafted features, such as cyclic-moment-based features extracted through expert domain knowledge [10667001]. A significant milestone was achieved in [o2016convolutional], where CNNs were introduced to learn features automatically, eliminating the dependence on expert-designed representations. This marked the advent of deep learning techniques in AMC.
@@
WTC is another challenging task that requires the model to accurately classify the type of wireless transmission based on the input IQ signal. Numerous studies have demonstrated significant progress in this area. For example, the authors in [FONTAINE2019101881] employed CNNs to automatically extract features from the input signal and used a feedforward neural network to classify the wireless technology. Data augmentation techniques were also incorporated, enabling the model to generalize well to unseen scenarios while maintaining high accuracy. In [9637487], a multiscale module was introduced to capture features at various resolutions. Additionally, a global average pooling layer was utilized to aggregate the extracted features, which reduced the number of parameters and accelerated the training process. Despite its simplified architecture, the model achieved strong classification performance and fast convergence.
The authors in [10.1016/j.vehcom.2022.100563] proposed a comprehensive WTC system designed to identify and analyze a wide range of wireless technologies coexisting within the 5.9 GHz band. The system specifically focused on LTE, Wi-Fi, 5G NR, C-V2X PC5, and ITS-G5. To ensure accurate identification, the study adopted a short time resolution window based on the minimum frame duration of the target technologies. Unlike most IQ-based approaches, MFBLN [10555427] leveraged alternative features such as received signal strength indicator (RSSI), fast fourier transform (FFT), and spectrogram representations to extract multi-dimensional signal characteristics. This allowed the model to capture richer information from the input, thereby enhancing classification accuracy and robustness across diverse wireless environments.
SS is a fundamental task in cognitive radio networks, referring to the process of monitoring and analyzing the RF spectrum to determine its occupancy status [10598350]. The primary goal of SS is to identify spectrum holes, which are unused or underutilized frequency bands in a specific geographical area. By detecting these holes, cognitive radios can dynamically access the available spectrum without interfering with primary users (PUs). SS is generally categorized into two types, namely narrowband SS and wideband SS. Narrowband SS analyzes one frequency band at a time, while wideband SS simultaneously evaluates multiple bands to identify vacant channels suitable for transmission [muzaffar2024review]. Common narrowband SS techniques include energy detection, cyclostationary detection, and matched filter detection [5723805]. Among these, energy detection is widely used due to its simplicity. It determines spectrum occupancy by comparing the measured signal energy against a predefined threshold.
@@ II-B Foundation Models
Foundation models are large-scale machine learning models that are pre-trained on vast amounts of data to learn general-purpose representations. The models serve as a foundational basis for a wide range of downstream tasks, making them highly versatile and adaptable across various domains. Foundation models are typically developed using deep learning techniques and are characterized by their massive scale, extensive training data, and strong transferability [10834497, 10173618]. The authors in [vaswani2017attention] introduced the attention mechanism and the Transformer architecture, which serve as the cornerstone of foundation models. BERT [devlin-etal-2019-bert] pioneered the era of foundation models. By introducing a bidirectional Transformer-based architecture and innovative pre-training strategies such as masked language modeling (MLM) and next sentence prediction (NSP), BERT achieved state-of-the-art performance on a wide range of nature language processing (NLP) tasks.
The BERT-like prediction loss is employed in HuBERT [10.1109/TASLP.2021.3122291] for self-supervised speech representation learning, leading to the development of a speech foundation model. The authors in [10.5555/3600270.3602542] present a foundational model that reformulates and unifies four core vision tasks, namely object detection, instance segmentation, keypoint prediction, and image captioning, into a single pixel-to-sequence framework. By framing multiple vision tasks within a shared pixel-to-sequence paradigm, this work demonstrated the potential of foundation models to generalize across tasks, reducing the need for specialized architectures and promoting greater flexibility in multimodal applications. SpectralGPT [10490262] represents a pioneering effort in creating a universal foundation model tailored for remote sensing (RS) applications. Trained on a large-scale dataset comprising one million spectral RS images, SpectralGPT achieved a parameter size exceeding 600 million. Evaluations across four downstream tasks reveal substantial performance improvements, highlighting the potential of foundation models to tackle a wide range of RS tasks.
WirelessGPT [yang2025wirelessgptgenerativepretrainedmultitask], a pioneering foundation model tailored for multi-task learning in ISAC, features an initial parameter size of approximately 80 million. Pre-trained on a large-scale wireless channel dataset, it delivers significant advancements in critical wireless tasks such as channel estimation, channel prediction, human activity recognition, and wireless reconstruction. WirelessGPT demonstrates that foundation models can achieve impressive performance improvements over conventional methods and smaller artificial intelligence (AI) models while requiring fewer labeled data, thereby enhancing efficiency and scalability in ISAC applications.
@@ III Our Proposed Model
$\Re(x)$ The real part of the signal.
$\Im(x)$ The imaginary part of the signal.
$\mathbf{A}$ The amplitude component.
$\boldsymbol{\varTheta}$ The phase component.
RNRN The residual connection and normalization operation.
$\text{concat}(\cdot)$ The concatenation operation.
@@
NN The number of signal symbols.
@@ III-A Data Collection And Processing Stage
In the data collection and processing stage, a diverse set of spectrum data is gathered from multiple sources, including publicly available open-source datasets and signals collected in real-world for this study, to facilitate comprehensive training of SpectrumFM. The open-source datasets utilized in this study include the RML2018.01A dataset [8267032] and the technology recognition (TechRec) dataset [^2] https://ieee-dataport.org/documents/iq-signals-captured-lte-wifi-and-dvb-t. The RML2018.01A dataset comprises 24 distinct modulation types, including OOK, 4-ASK, 8-ASK, BPSK, QPSK, 8-PSK, 16-PSK, 32-PSK, 16-APSK, 32-APSK, 64-APSK, 128-APSK, 16-QAM, 32-QAM, 64-QAM, 128-QAM, 256-QAM, AM-SSB-WC, AM-SSB-SC, AM-DSB-WC, AM-DSB-SC, FM, GMSK, and OQPSK. This dataset spans a broad range of SNRs, ranging from -20 dB to 30 dB in increments of 2 dB, providing a comprehensive representation of diverse signal conditions. This dataset mainly reflects typical wireless communication scenarios across different modulation schemes under various channel impairments, such as fading, frequency offsets, and noise. The TechRec dataset consists of over-the-air IQ samples collected from six distinct locations in Gent, Belgium, including UZ, Reep, Rabot, Merelbeke, iGent, and Gentbrugge. Each location exhibits unique signal characteristics due to variations in signal strength, noise levels, and physical propagation conditions. The SNR in the dataset ranges from -6 dB to 12 dB, with increments of 2 dB. Specifically, IQ samples collected from Merelbeke, iGent, and Rabot are leveraged for pre-training, while data from UZ, Reep, and Gentbrugge are exploited for fine-tuning and testing. This dataset reflects real-world urban communication scenarios, capturing the effects of multipath propagation, background interference, and heterogeneous spectrum occupancy across different city environments. In addition to the publicly available datasets, signals are also collected by using a Ceyear 1435B-V RF signal generator for transmission and a SAM 60 MK2 receiver for reception. This self-collected dataset covers both indoor scenarios, such as office-like multipath-rich environments, and outdoor scenarios, such as open-space propagation conditions, thereby complementing the public datasets with more diverse real-world signal characteristics.
TABLE II: Overview of The Pretraining Datasets Used in This Study.
Dataset # Samples Sample Shape Approx. Contribution
RML2018.01A 2,555,904 1024 $\times$ 2 52.1%
TechRec 9,747,800 128 $\times$ 2 24.9%
Self-collected $\sim$9,000,000 128 $\times$ 2 23.0%
The total size of the pre-training dataset amounts to approximately 25 GB. A detailed breakdown of the datasets is shown in TABLE II. The received signal $H$ is formulated as
@@
To facilitate normalization, we first convert the IQ signals into amplitude-phase (AP) signals, since the value ranges of amplitude and phase components are better suited for standardization, specifically $\pi,\pi]$ for the phase component and $0,+\infty]$ for the amplitude component. Mathematically, this transformation for the IQ signals to AP signals can be expressed as
@@ III-B SpectrumFM Encoder
where $\mathbf{x}_{\text{proj}}\in\mathbb{R}^{N\times d}$ is the projected signal, $\mathbf{W}_{\text{proj}}\in\mathbb{R}^{d\times 2\times 1}$ is the convolution kernel, $d$ is the latent dimension, and $p$ represents the position index, ranging from 0 to N1N-1. To capture the sequential order in a signal sequence, positional encoding is added to $\mathbf{x}_{\text{proj}}$. The positional encoding is given as
@@
where $i$ is the index of the dimension, ranging from 0 to d1d-1. The signal with the positional encoding is given as
@@
where concat()($\cdot$) is the concatenation operator, and $\mathbf{W}^{o}\in\mathbb{R}^{d\times d}$ is the weight matrix.
@@ III-C1 The Masked Reconstruction Task
In the masked reconstruction task, the objective is to mask a portion of the original AP sequence $\mathbf{x}_{\text{ap}}^{\text{norm}}$ and train the model to reconstruct the masked parts. The task facilitates learning general representations in an unsupervised manner by training the model to infer and reconstruct masked portions of the input data, thereby enhancing its robustness and generalization capabilities. Specifically, given the sequence $\mathbf{x}_{\text{ap}}^{\text{norm}}$, we first randomly mask a subset of the sequence according to a predefined masking ratio $r$. The binary mask vector $\mathbf{m}$ is defined as $\mathbf{m}\sim\text{Bernoulli}(1-r)$, which consists of elements $\left\{m_{1},m_{2},\cdots,m_{N}\right\}$. Here $m_{p}=0$ if the $p$-th symbol is masked out, and $m_{p}=1$ otherwise. The masked sequence is then given as
@@
where $\mathbf{M}$ assigns extremely small values (e.g., -$\infty$) to the positions that are masked out, thereby ensuring that the positions do not contribute to the computed attention weights. After obtaining $\mathbf{x}_{\text{{latent}}}^{(L)}$, a lightweight decoder is applied to the latent representation to reconstruct the original input sequence, given as
@@ III-C2 Next-slot Signal Prediction Task
The next-slot signal prediction task aims to predict the signal symbol at the $N$-th time slot based on the historical signal observations from the previous N1N-1 slots. The task facilitates SpectrumFM to understand the temporal characteristics of signals. Specifically, $\mathbf{x}_{\text{ap}}^{\text{norm}}[:N-1]$ is input to the SpectrumFM encoders to obtain the latent representation $\mathbf{x}_{\text{{latent}}}^{(L)}$, and then an long short-term memory (LSTM) layer is leveraged to aggregate the latent representation, and finally a linear layer is exploited to predict the $N$-th signal symbol. Mathematically, the predicted signal symbol is given as
@@ III-D Fine-tuning Stage
III-E Theoretical Computational Complexity
To further clarify the computational footprint of SpectrumFM, we provide a theoretical analysis of the per-layer complexity. Each encoder layer consists of a multi-head self-attention (MHSA) module, two feed-forward (FFN) modules, and a convolution (Conv) module. Let $N$ denote the input sequence length, $d$ be the latent dimension, $h$ be the number of attention heads, $d_{h}=d/h$ be the dimension per head, $d_{\text{feed}}$ be the feed-forward hidden dimension, $k$ be the kernel size of the depthwise convolution, and $L$ be the number of layers. The complexity of the MHSA module can be expressed as $\mathcal{O}_{\text{MHSA}}=4Nd^{2}+2hN^{2}d_{h}$, where the first term accounts for the linear projections of queries, keys, and values, and the second term corresponds to the scaled dot-product attention. Each feed-forward module contributes $\mathcal{O}_{\text{FFN}}=2Ndd_{\text{feed}}$, and the convolution module contributes $\mathcal{O}_{\text{Conv}}=Ndk$. Therefore, the total complexity per layer is approximately $\mathcal{O}_{\text{layer}}=\mathcal{O}_{\text{MHSA}}+2\,\mathcal{O}_{\text{FF}}+\mathcal{O}_{\text{Conv}}$, and for a $L$-layer encoder, the overall forward pass complexity is $\mathcal{O}_{\text{encoder}}=L\mathcal{O}_{\text{layer}}$. Substituting our model parameters ($N=128$, $d=256$, $h=4$, $d_{\text{feed}}=512$, $k=3$, $L=16$) yields a theoretical per-sample FLOPs estimate of approximately $2.03\times 10^{9}$ FLOPs. This analysis provides a clear understanding of the computational demands associated with SpectrumFM, facilitating informed decisions regarding its deployment in various applications.